A certain traveling wave has frequency of 200hz. A wave length of 2m and an amplitude of 0.04m. calculate the displacement (Y) at a point (x) 0,3 from the origin at a time 0.01sec after zero displacement at that point (note, the wave is traveling in opposite direction).


1
Expert's answer
2022-04-27T13:34:46-0400

"s(x,t)=-A\cos(2\pi\nu t-\frac{2\pi x}{\lambda}),"

"s(0.3,~0.01)=-0.024~m."